# Generalized Homogeneous Derivations on Graded Rings

### v5

@@ 6. Generalized Homogeneous Derivations on Graded Modules
Let $R$ be a $G$-graded ring and $M$ a $G$-graded $R$-module. An additive mapping $F_{M}:M\to M$ is a generalized homogeneous derivation if there exists a homogeneous derivation $d:R\to R$ such that
  1. (i)
@@
    If $(F_{M},d)_{h,M},(F_{N},d)_{h,N}\in p\mathfrak{Der}^{gh}_{G}(R,M),p\mathfrak{Der}^{gh}_{G}(R,N)$ have the same associated derivation $d$, then
@@
For degree preservation, let $(m_{i})_{i\in I}$ be homogeneous of degree $\tau\in G$ in the direct sum. Then each nonzero $m_{i}$ is homogeneous of degree $\tau$, and since each $F_{M_{i}}$ preserves degrees, $F_{M_{i}}(m_{i})$ is either zero or homogeneous of degree $\tau$. Hence $(F_{M_{i}}(m_{i}))_{i\in I}$ is homogeneous of degree $\tau$, so $F_{\bigoplus M_{i}}$ preserves the grading.
(2) For $r\in R$, $m\in M_{\tau}$, and $n\in N_{\sigma}$, we have
$F_{M\otimes N}(r(m\otimes n))$ $=F_{M\otimes N}(rm\otimes n)$
$=F_{M}(rm)\otimes n+rm\otimes F_{N}(n)$
@@
$=d(r)m\otimes n+rF_{M}(m)\otimes n+rm\otimes F_{N}(n)$
$=d(r)(m\otimes n)+r(F_{M}(m)\otimes n+m\otimes F_{N}(n))$
$=d(r)(m\otimes n)+rF_{M\otimes N}(m\otimes n).$
For degree preservation, if $m\in M_{\tau}$ and $n\in N_{\sigma}$, then $m\otimes n\in(M\otimes_{R}N)_{\tau\sigma}$. Since $F_{M}(m)\in M_{\tau}$ et $F_{N}(n)\in N_{\sigma}$, we have
$F_{M\otimes N}(m\otimes n)=F_{M}(m)\otimes n+m\otimes F_{N}(n)\in(M\otimes_{R}N)_{\tau\sigma}.$
Thus $F_{M\otimes N}$ preserves the grading. ∎
@@
A graded $R$-module homomorphism $\phi:M\to N$ is a gr-generalized homomorphism if
$\phi\circ F_{M}=F_{N}\circ\phi$
for $(F_{M},d)_{h,M}\in p\mathfrak{Der}^{gh}_{G}(R,M)$ and $(F_{N},d)_{h,N}\in p\mathfrak{Der}^{gh}_{G}(R,N)$.
@@
Let $\{M_{i}\}_{i\in I}$ be a finite family of graded $R$-modules with direct sum $M=\bigoplus_{i\in I}M_{i}$. If each $M_{i}$ admits a gr-generalized derivation $(F_{M_{i}},d)_{h,M_{i}}\in p\mathfrak{Der}^{gh}_{G}(R,M_{i})$ with the same associated derivation $d$, then the canonical projection maps $\pi_{j}:M\to M_{j}$ are gr-generalized homomorphisms with respect to the gr-generalized derivations $(F_{M},d)_{h,M}$ on $M$ and $(F_{M_{j}},d)_{h,M_{j}}$ on $M_{j}$.
@@
such that for any $(F_{M},d)_{h,M}\in p\mathfrak{Der}^{gh}_{G}(R,M)$ with $F_{M}(\ker\phi)\subseteq\ker\phi$, the induced map is $(F_{N},d)_{h,N}=\phi_{*}((F_{M},d)_{h,M})$.
@@
Since $\phi$ is surjective, for each $n\in N$ there exists $m\in M$ with $\phi(m)=n$. Define $F_{N}:N\to N$ by
$F_{N}(n)=\phi(F_{M}(m)),$
where $m$ is any preimage of $n$. To see that $F_{N}$ is well defined, suppose $\phi(m_{1})=\phi(m_{2})=n$. Then $m_{1}-m_{2}\in\ker\phi$, and by hypothesis
$F_{M}(m_{1}-m_{2})\in\ker\phi.$
Hence $\phi(F_{M}(m_{1}))=\phi(F_{M}(m_{2}))$. For the gr-generalized derivation property, let $r\in R$ and $n\in N$, and choose $m\in M$ with $\phi(m)=n$. Then
$F_{N}(rn)$ $=F_{N}(\phi(rm))$
$=\phi(F_{M}(rm))$
$=\phi(d(r)m+rF_{M}(m))$
$=d(r)\phi(m)+r\phi(F_{M}(m))$
$=d(r)n+rF_{N}(n).$
For homogeneity, if $n\in N_{\tau}$, then, since $\phi$ is graded, we can choose $m\in M_{\tau}$ with $\phi(m)=n$. As $F_{M}$ preserves degrees, $F_{M}(m)\in M_{\tau}$, and thus
$F_{N}(n)=\phi(F_{M}(m))\in\phi(M_{\tau})\subseteq N_{\tau}.$
The $Z(R)\cap R_{e}$-linearity of $\phi_{*}$ follows from the linearity of $\phi$ and the module structure on $p\mathfrak{Der}^{gh}_{G}(R,M)$. ∎
@@
For graded isomorphisms $\phi:M\to N$, the induced map
$\phi_{*}:p\mathfrak{Der}^{gh}_{G}(R,M)\to p\mathfrak{Der}^{gh}_{G}(R,N)$
is a $Z(R)\cap R_{e}$-module isomorphism, with inverse
$\psi_{*}((F_{N},d)_{h,N})=(\phi^{-1}\circ F_{N}\circ\phi,d)_{h,M}.$
@@
The canonical projection $\pi:M\to M/N$ is surjective, and since $F_{M}(N)\subseteq N$ by hypothesis, we have $F_{M}(\ker\pi)\subseteq\ker\pi$. By Proposition 6.3, there exists a well-defined $Z(R)\cap R_{e}$-linear map
$\pi_{*}:p\mathfrak{Der}^{gh}_{G}(R,M)\to p\mathfrak{Der}^{gh}_{G}(R,M/N)$
such that $\pi_{*}((F_{M},d)_{h,M})=(F_{M/N},d)_{h,M/N}$, where $F_{M/N}(\pi(m))=\pi(F_{M}(m))$ for all $m\in M$. Well-definedness of $F_{M/N}$ follows from $F_{M}(N)\subseteq N$. The gr-generalized derivation property is checked by verifying
$F_{M/N}(r\overline{m})=d(r)\overline{m}+rF_{M/N}(\overline{m})$
for all $r\in R$ and $\overline{m}\in M/N$. For homogeneity, if $\overline{m}\in(M/N)_{\tau}$, write $m=m_{\tau}+n$ with $m_{\tau}\in M_{\tau}$ and $n\in N$. Since $F_{M}$ preserves degrees and $F_{M}(N)\subseteq N$, we get
$F_{M/N}(\overline{m})=\pi(F_{M}(m_{\tau}))\in(M/N)_{\tau}.$
The commutativity relation $F_{M/N}\circ\pi=\pi\circ F_{M}$ makes $\pi$ a gr-generalized homomorphism. Uniqueness follows because any other gr-generalized derivation $F^{\prime}_{M/N}$ satisfying $F^{\prime}_{M/N}\circ\pi=\pi\circ F_{M}$ must coincide with $F_{M/N}$ on all cosets. ∎
@@
  2. (ii)
    Morphisms: Pairs $(\phi,\psi):(R,M,(F_{M},d)_{h,M})\to(S,N,(F_{N},e)_{h,N})$ where $\phi:R\to S$ is a graded ring homomorphism, $\psi:M\to N$ is $\phi$-semilinear, and the diagrams
@@
Composition is well defined: Let
$(\phi,\psi):(R,M,(F_{M},d)_{h,M})\to(S,N,(F_{N},e)_{h,N})$
and
$(\phi^{\prime},\psi^{\prime}):(S,N,(F_{N},e)_{h,N})\to(T,P,(F_{P},f)_{h,P})$
be morphisms in $\mathscr{M}^{gh}_{G}$. We must show that $(\phi^{\prime}\circ\phi,\psi^{\prime}\circ\psi)$ is again a morphism. First, $\psi^{\prime}\circ\psi$ is $(\phi^{\prime}\circ\phi)$-semilinear:
$(\psi^{\prime}\circ\psi)(rm)$ $=\psi^{\prime}(\psi(rm))$
$=\psi^{\prime}(\phi(r)\psi(m))$
$=\phi^{\prime}(\phi(r))\psi^{\prime}(\psi(m))$
$=(\phi^{\prime}\circ\phi)(r)(\psi^{\prime}\circ\psi)(m).$
Next, we check compatibility with the derivations and module maps:
$(\psi^{\prime}\circ\psi)\circ F_{M}=\psi^{\prime}\circ(F_{N}\circ\psi)=F_{P}\circ(\psi^{\prime}\circ\psi),$
and similarly
$(\phi^{\prime}\circ\phi)\circ d=f\circ(\phi^{\prime}\circ\phi).$
Identity morphisms: For any object $(R,M,(F_{M},d)_{h,M})$, the pair $(\text{id}_{R},\text{id}_{M})$ satisfies
$\text{id}_{M}(rm)=rm=\text{id}_{R}(r)\text{id}_{M}(m),\quad\text{id}_{M}\circ F_{M}=F_{M}=F_{M}\circ\text{id}_{M},\quad\text{id}_{R}\circ d=d=d\circ\text{id}_{R},$
and is therefore a morphism in $\mathscr{M}^{gh}_{G}$.
Associativity and identity laws: These follow directly from the associativity and identity properties of function composition. ∎

### v6

@@ 6. Generalized homogeneous derivations on graded modules
Let $R$ be a $G$-graded ring and $M$ a $G$-graded $R$-module. An additive mapping $F_{M}:M\to M$ is a generalized homogeneous derivation if there exists a homogeneous derivation $d:R\to R$ such that:
  1. (1)
@@
    If $(F_{M},d)_{h,M}$ and $(F_{N},d)_{h,N}$ have the same associated derivation $d$, then
@@
For degree preservation, if $(m_{i})_{i\in I}$ is homogeneous of degree $\tau$, then each nonzero $m_{i}\in M_{i,\tau}$, and $F_{M_{i}}(m_{i})\in M_{i,\tau}$ by hypothesis.
(2) For $r\in R$, $m\in M_{\tau}$, $n\in N_{\sigma}$:
$F_{M\otimes N}(r(m\otimes n))$ $=F_{M}(rm)\otimes n+rm\otimes F_{N}(n)$
@@
$=d(r)(m\otimes n)+r(F_{M}(m)\otimes n+m\otimes F_{N}(n))=d(r)(m\otimes n)+rF_{M\otimes N}(m\otimes n).$
Degree preservation follows from $F_{M}(m)\in M_{\tau}$, $F_{N}(n)\in N_{\sigma}$, and $m\otimes n\in(M\otimes_{R}N)_{\tau\sigma}$. ∎
@@
A graded $R$-module homomorphism $\phi:M\to N$ is a gr-generalized homomorphism if $\phi\circ F_{M}=F_{N}\circ\phi$ for $(F_{M},d)_{h,M}\in p\mathfrak{Der}^{gh}_{G}(R,M)$ and $(F_{N},d)_{h,N}\in p\mathfrak{Der}^{gh}_{G}(R,N)$.
@@
Let $\{M_{i}\}_{i\in I}$ be a finite family of graded $R$-modules with direct sum $M=\bigoplus_{i\in I}M_{i}$. If each $M_{i}$ admits a gr-generalized derivation $(F_{M_{i}},d)_{h,M_{i}}$ with the same associated derivation $d$, then the canonical projection maps $\pi_{j}:M\to M_{j}$ are gr-generalized homomorphisms.
@@
such that for any $(F_{M},d)_{h,M}$ with $F_{M}(\ker\phi)\subseteq\ker\phi$, the induced map is $(F_{N},d)_{h,N}=\phi_{*}((F_{M},d)_{h,M})$.
@@
Since $\phi$ is surjective, for each $n\in N$ choose $m\in M$ with $\phi(m)=n$ and define $F_{N}(n)=\phi(F_{M}(m))$. If $\phi(m_{1})=\phi(m_{2})=n$, then $m_{1}-m_{2}\in\ker\phi$ and $F_{M}(m_{1}-m_{2})\in\ker\phi$, so $\phi(F_{M}(m_{1}))=\phi(F_{M}(m_{2}))$, giving well-definedness. For any $r\in R$ and $n\in N$, choosing $m$ with $\phi(m)=n$:
$F_{N}(rn)=\phi(F_{M}(rm))=\phi(d(r)m+rF_{M}(m))=d(r)n+rF_{N}(n).$
For homogeneity, if $n\in N_{\tau}$ choose $m\in M_{\tau}$ with $\phi(m)=n$; then $F_{M}(m)\in M_{\tau}$ and $F_{N}(n)=\phi(F_{M}(m))\in N_{\tau}$. The $Z(R)\cap R_{e}$-linearity of $\phi_{*}$ follows from the linearity of $\phi$ and the module structure on $p\mathfrak{Der}^{gh}_{G}(R,M)$. ∎
@@
For graded isomorphisms $\phi:M\to N$, the induced map $\phi_{*}:p\mathfrak{Der}^{gh}_{G}(R,M)\to p\mathfrak{Der}^{gh}_{G}(R,N)$ is a $Z(R)\cap R_{e}$-module isomorphism, with inverse $\psi_{*}((F_{N},d)_{h,N})=(\phi^{-1}\circ F_{N}\circ\phi,d)_{h,M}$.
@@
The canonical projection $\pi:M\to M/N$ is surjective with $F_{M}(\ker\pi)\subseteq\ker\pi$ by hypothesis. By Proposition 6.10, $\pi_{*}((F_{M},d)_{h,M})=(F_{M/N},d)_{h,M/N}$ where $F_{M/N}(\pi(m))=\pi(F_{M}(m))$. Well-definedness follows from $F_{M}(N)\subseteq N$. The gr-generalized derivation property and homogeneity of $F_{M/N}$ are verified directly. The commutativity relation $F_{M/N}\circ\pi=\pi\circ F_{M}$ makes $\pi$ a gr-generalized homomorphism, and uniqueness follows from the universal property. ∎
@@
  2. (2)
    Morphisms: Pairs $(\phi,\psi):(R,M,(F_{M},d)_{h,M})\to(S,N,(F_{N},e)_{h,N})$ where $\phi:R\to S$ is a graded ring homomorphism, $\psi:M\to N$ is $\phi$-semilinear, and the following diagrams commute:
@@
Composition. Let $(\phi,\psi)$ and $(\phi^{\prime},\psi^{\prime})$ be composable morphisms. Then $\psi^{\prime}\circ\psi$ is $(\phi^{\prime}\circ\phi)$-semilinear:
$(\psi^{\prime}\circ\psi)(rm)=\psi^{\prime}(\phi(r)\psi(m))=(\phi^{\prime}\circ\phi)(r)(\psi^{\prime}\circ\psi)(m).$
Compatibility with derivations and module maps:
$(\psi^{\prime}\circ\psi)\circ F_{M}=\psi^{\prime}\circ(F_{N}\circ\psi)=F_{P}\circ(\psi^{\prime}\circ\psi),\quad(\phi^{\prime}\circ\phi)\circ d=f\circ(\phi^{\prime}\circ\phi).$
Identity morphisms. For $(R,M,(F_{M},d)_{h,M})$, the pair $(\mathrm{id}_{R},\mathrm{id}_{M})$ satisfies all required conditions.
Associativity and identity laws follow from the corresponding properties of function composition. ∎